Ran diagnostics across the Chorus Array; signal drift dropped to green.
Collected fresh data for the Chorus Array; calibrations read 66% stable. Requesting cross-checks on recording controlled anomalies; please attach logs before the next sync window.
Action queue:
- Invite collaborators
- Document fresh findings
- Note replication steps
計画を整えて、次の波へ。
Let’s keep the worm cheering us on, not cleaning up after us.